# Modewise Additive Factor Model for Matrix Time Series

### v1

@@ 4.2 Statistical Inference for Factor Loadings
In factor models, each row of the loading matrix represents how a particular unit (e.g. a country or an asset) loads onto the latent factors. For example, in finance, asset pricing models posit that each stock’s return depends linearly on a set of common factors, with the factor loadings capturing the asset’s risk exposures or beta coefficients, to systematic market factors (ROSS1976341; connor1986performance; lettau2022high; giglio2022factor; conlon2025asset). In macroeconomics, a loading row may measure a country’s sensitivity to global business cycle shocks (bai2003inferential; houssa2008monetary; forni2021policy). Inference on these row-specific loadings is crucial for testing hypotheses about unit heterogeneity (e.g., which countries significantly load on a global factor) or about the stability of the factor structure over time (e.g., whether certain assets cease to load on a market factor).
@@
Theorem 4.4 establishes the asymptotic normality of each row of the estimated factor loading matrices, with means aligned to the true values via rotations $\boldsymbol{R}_{\boldsymbol{A}},\boldsymbol{R}_{\boldsymbol{B}}$. The asymptotic variances involve the column-wise covariance matrices of the projected factors ($\boldsymbol{\Sigma}_{\boldsymbol{F}}$ and $\boldsymbol{\Sigma}_{\boldsymbol{G}}$), and the covariance matrices of the signal-by-noise cross-products ($\boldsymbol{\Sigma}_{\boldsymbol{A},i}$ and $\boldsymbol{\Sigma}_{\boldsymbol{B},j}$), analogous to statistical inference in conventional factor models. However, our inferential framework relies on a fine-grained analysis of the covariance matrix of the projected data. Specifically, we employ spectral representation tools from xia2021normal to achieve a second-order accurate characterization of matrix SVD, enabling precise control of the leading term and sharp bounds on higher-order terms. This differs fundamentally from techniques used in vector factor models (bai2003inferential) and Tucker factor models (chen2023statistical; yu2022projected), which adopt the approach of bai2003inferential. Our framework offers two advantages: first, we establish explicit convergence rates for asymptotic normality, which previous methods cannot provide; second, the technique in bai2003inferential relies on a manually designed rotation matrix for error decomposition, which does not extend to our setting due to the modewise additive structure and high dimensionality of the factor processes. Our approach can be generalized to inference in other matrix and tensor factor models, such as CP factor models (chang2023modelling; chang2024identification).
@@
In the strong factor model case with $\delta_{0}=\delta_{1}=0$ and $d_{1}\asymp d_{2}$, we have $\Omega_{\boldsymbol{A}}\asymp\Omega_{\boldsymbol{B}}\asymp\sqrt{n}d_{1}^{-1}$. Thus, asymptotic normality in Theorem 4.4 requires $n\ll d_{1}^{2}$. This condition reflects a fundamental phase transition in high-dimensional factor models, sharing the same spirit as Theorem 1 of bai2003inferential and Theorem 2 of chen2023statistical. When $n\ll d_{1}^{2}$, the first-order signal-by-noise cross-product dominates the perturbation $\widehat{\boldsymbol{U}}_{\boldsymbol{A}}-\boldsymbol{U}_{\boldsymbol{A}}\boldsymbol{R}_{\boldsymbol{A}}$, yielding the asymptotic normality stated in the theorem. Otherwise, the first-order noise-by-noise cross-product becomes the dominant contributor, introducing non-negligible bias that requires more sophisticated distributional characterization involving the error covariance structure. The latter case parallels traditional PCA, where debiasing is often necessary to achieve asymptotic normality for linear combinations of principal components (koltchinskii2020efficient). Notably, if the entries of noise process $\boldsymbol{E}_{t}$ are i.i.d., the condition $n\ll d_{1}^{2}$ can be removed, as the population noise covariance does not affect the singular space of the covariance of the projected data.
@@ 4.3 Data-Driven Statistical Inference
Although statistical inference for loading matrices has been studied in vector (bai2003inferential), Tucker (chen2023statistical), and CP (chen2026estimation) factor models, practical procedures for constructing confidence intervals remain underexplored, where covariance matrix estimation is crucial, particularly under weak factor settings. In this subsection, we develop a data-driven inference approach based on plug-in estimators. We first estimate the projected factor covariance matrices, which capture the row-wise covariances of the projected factors $\boldsymbol{U}_{\boldsymbol{B}\perp}\boldsymbol{U}_{\boldsymbol{B}\perp}^{\top}\widetilde{\mathbf{\boldsymbol{F}}}_{t}$ and $\boldsymbol{U}_{\boldsymbol{A}\perp}\boldsymbol{U}_{\boldsymbol{A}\perp}^{\top}\widetilde{\mathbf{\boldsymbol{G}}}_{t}$. Using the estimated factors in (5), the plug-in estimators for the projected factor covariance matrices in (14) are
@@ 7 Real Data Analysis
To remove scale differences across indicators, we apply pooled standardization: for each indicator $j\in[d_{2}]$, we compute the pooled mean $\mu_{j}$ and standard deviation $\sigma_{j}$ across all country-quarter observations and set $X_{t,ij}=(Y_{t,ij}-\mu_{j})/\sigma_{j}$. This follows standard practice in macroeconomic panel studies (ottinger2025history). All estimation and forecasting use the standardized series $\{\boldsymbol{X}_{t}\}_{t=1}^{n}$.
Competing methods. We compare MAFM using MINE initialization and COMPAS refinement against four alternative approaches: (1) TwDFM (yuan2023twoway), which decomposes matrix time series into separate row and column components with autoregressive dynamics via two-step quasi-maximum likelihood estimation; (2) Tucker factor model (han2024tensor) with cross-covariance based estimation; (3) CP factor model (chen2026estimation) with cross-covariance based estimation; and (4) VecFactor (bai2003inferential), a vectorized baseline that flattens matrix observations into vectors before applying standard factor analysis with cross-covariance estimation. For fair comparison, all methods use AR models to forecast factors, which are then used to reconstruct predicted observations.

### v2

@@ 4.2 Statistical Inference for Factor Loadings
In factor models, each row of the loading matrix represents how a particular unit (e.g. a country or an asset) loads onto the latent factors. For example, in finance, asset pricing models posit that each stock’s return depends linearly on a set of common factors, with the factor loadings capturing the asset’s risk exposures or beta coefficients, to systematic market factors (Ross, 1976; Connor and Korajczyk, 1986; Lettau, 2022; Giglio et al., 2022; Conlon et al., 2025). In macroeconomics, a loading row may measure a country’s sensitivity to global business cycle shocks (Bai, 2003; Houssa, 2008; Forni and Gambetti, 2021). Inference on these row-specific loadings is crucial for testing hypotheses about unit heterogeneity (e.g., which countries significantly load on a global factor) or about the stability of the factor structure over time (e.g., whether certain assets cease to load on a market factor).
@@
Theorem 4.4 establishes the asymptotic normality of each row of the estimated factor loading matrices, with means aligned to the true values via rotations $\boldsymbol{R}_{\boldsymbol{A}},\boldsymbol{R}_{\boldsymbol{B}}$. The asymptotic variances involve the column-wise covariance matrices of the projected factors ($\boldsymbol{\Sigma}_{\boldsymbol{F}}$ and $\boldsymbol{\Sigma}_{\boldsymbol{G}}$), and the covariance matrices of the signal-by-noise cross-products ($\boldsymbol{\Sigma}_{\boldsymbol{A},i}$ and $\boldsymbol{\Sigma}_{\boldsymbol{B},j}$), analogous to statistical inference in conventional factor models. However, our inferential framework relies on a fine-grained analysis of the covariance matrix of the projected data. Specifically, we employ spectral representation tools from Xia (2021) to achieve a second-order accurate characterization of matrix SVD, enabling precise control of the leading term and sharp bounds on higher-order terms. This differs fundamentally from techniques used in vector factor models (Bai, 2003) and Tucker factor models (Chen and Fan, 2023; Yu et al., 2022), which adopt the approach of Bai (2003). Our framework offers two advantages: first, we establish explicit convergence rates for asymptotic normality, which previous methods cannot provide; second, the technique in Bai (2003) relies on a manually designed rotation matrix for error decomposition, which does not extend to our setting due to the modewise additive structure and high dimensionality of the factor processes. Our approach can be generalized to inference in other matrix and tensor factor models, such as CP factor models (Chang et al., 2023, 2024).
@@
In the strong factor model case with $\delta_{0}=\delta_{1}=0$ and $d_{1}\asymp d_{2}$, we have $\Omega_{\boldsymbol{A}}\asymp\Omega_{\boldsymbol{B}}\asymp\sqrt{n}d_{1}^{-1}$. Thus, asymptotic normality in Theorem 4.4 requires $n\ll d_{1}^{2}$. This condition reflects a fundamental phase transition in high-dimensional factor models, sharing the same spirit as Theorem 1 of Bai (2003) and Theorem 2 of Chen and Fan (2023). When $n\ll d_{1}^{2}$, the first-order signal-by-noise cross-product dominates the perturbation $\widehat{\boldsymbol{U}}_{\boldsymbol{A}}-\boldsymbol{U}_{\boldsymbol{A}}\boldsymbol{R}_{\boldsymbol{A}}$, yielding the asymptotic normality stated in the theorem. Otherwise, the first-order noise-by-noise cross-product becomes the dominant contributor, introducing non-negligible bias that requires more sophisticated distributional characterization involving the error covariance structure. The latter case parallels traditional PCA, where debiasing is often necessary to achieve asymptotic normality for linear combinations of principal components (Koltchinskii et al., 2020). Notably, if the entries of noise process $\boldsymbol{E}_{t}$ are i.i.d., the condition $n\ll d_{1}^{2}$ can be removed, as the population noise covariance does not affect the singular space of the covariance of the projected data.
@@ 4.3 Data-Driven Statistical Inference
Although statistical inference for loading matrices has been studied in vector (Bai, 2003), Tucker (Chen and Fan, 2023), and CP (Chen et al., 2026) factor models, practical procedures for constructing confidence intervals remain underexplored, where covariance matrix estimation is crucial, particularly under weak factor settings. In this subsection, we develop a data-driven inference approach based on plug-in estimators. We first estimate the projected factor covariance matrices, which capture the row-wise covariances of the projected factors $\boldsymbol{U}_{\boldsymbol{B}\perp}\boldsymbol{U}_{\boldsymbol{B}\perp}^{\top}\widetilde{\mathbf{\boldsymbol{F}}}_{t}$ and $\boldsymbol{U}_{\boldsymbol{A}\perp}\boldsymbol{U}_{\boldsymbol{A}\perp}^{\top}\widetilde{\mathbf{\boldsymbol{G}}}_{t}$. Using the estimated factors in (5), the plug-in estimators for the projected factor covariance matrices in (14) are
@@ 7 Real Data Analysis
To remove scale differences across indicators, we apply pooled standardization: for each indicator $j\in[d_{2}]$, we compute the pooled mean $\mu_{j}$ and standard deviation $\sigma_{j}$ across all country-quarter observations and set $X_{t,ij}=(Y_{t,ij}-\mu_{j})/\sigma_{j}$. This follows standard practice in macroeconomic panel studies (Ottinger and Voigtländer, 2025). All estimation and forecasting use the standardized series $\{\boldsymbol{X}_{t}\}_{t=1}^{n}$.
Competing methods. We compare MAFM using MINE initialization and COMPAS refinement against four alternative approaches: (1) TwDFM (Yuan et al., 2023), which decomposes matrix time series into separate row and column components with autoregressive dynamics via two-step quasi-maximum likelihood estimation; (2) Tucker factor model (Han et al., 2024a) with cross-covariance based estimation; (3) CP factor model (Chen et al., 2026) with cross-covariance based estimation; and (4) VecFactor (Bai, 2003), a vectorized baseline that flattens matrix observations into vectors before applying standard factor analysis with cross-covariance estimation. For fair comparison, all methods use AR models to forecast factors, which are then used to reconstruct predicted observations.